### v1

@@ 4.2 Converting Templates to Descriptions
MolT5-Small 29.83 3.48 3.12 15.13 4.18 4.29 12.42 4.85 3.27 46.77 0.57 0.56 45.00 4.32 4.36 0.00 0.00 0.00
MolT5-Base 35.36 5.18 4.69 14.58 4.84 4.97 16.08 5.82 3.36 63.94 5.01 5.18 46.85 5.05 5.27 0.00 0.00 0.00
MolT5-Large 33.32 7.72 6.95 15.27 7.94 7.82 16.96 10.90 7.39 62.77 5.99 6.27 38.29 6.06 6.31 0.00 0.00 0.00
Meditron-7B 25.27 11.56 12.04 23.86 14.91 16.22 26.51 16.48 17.49 29.54 7.52 7.07 21.18 7.35 7.40 12.35 0.29 0.56
@@
MolT5-Small 0.00 0.00 0.00 0.00 0.00 0.00 24.85 9.69 6.54 0.00 0.00 0.00 3.42 0.43 0.09 1.96 0.00 0.00
MolT5-Base 0.00 0.00 0.00 67.45 8.51 8.84 28.00 11.51 6.52 4.17 0.12 0.20 2.20 0.58 0.11 9.70 0.23 0.15
MolT5-Large 0.00 0.00 0.00 69.42 10.29 10.85 15.77 12.28 8.16 18.14 9.52 6.62 8.90 2.28 1.13 4.32 1.16 0.61
Meditron-7B 0.00 0.00 0.00 48.79 11.75 11.05 29.10 20.64 20.64 23.93 12.33 14.34 35.69 19.91 22.65 14.79 9.34 8.98
@@
MolT5-Small 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 55.49 1.99 1.70 87.44 50.08 49.94 71.86 21.03 24.27
MolT5-Base 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 58.90 2.25 1.80 94.61 55.16 59.18 45.06 25.49 24.54
@@ 6 Benchmarks
Overall, the dataset proves to be fairly challenging for these naively finetuned models. On captioning, Meditron-7B achieves a maximum overall F-1 score of 12.04 for property identification (Table 3). However, overall it has a much higher precision than recall, indicating the model only labels a molecule with a certain property when having higher confidence. Certain classes of molecules, such as X-icides, are never identified (Table 4). Other classes, such as toxins or electricity, show emergent behavior as model size scales. Interestingly, the models appear to be fairly capable at linking molecules to certain diseases or cancers. We find that, likely due to poor performance on individual properties, only the largest model succeeds on predicting held-out combos, and with poor results. Additionally, we find that the Text2Mol metric, as trained on ChEBI-20, shows poor domain transfer to L+M-24. For the final workshop evaluation, we plan to train an updated Text2Mol model on L+M-24.

### v2

@@ 4.2 Converting Templates to Descriptions
MolT5-Small 84.83 8.24 7.88 85.13 23.23 23.33 62.42 4.85 3.27 96.77 0.57 0.56 95.00 4.32 4.36 0.00 0.00 0.00
MolT5-Base 64.11 9.94 9.46 79.58 23.89 24.02 16.08 5.82 3.36 63.94 5.01 5.18 96.85 5.05 5.27 0.00 0.00 0.00
MolT5-Large 59.57 12.49 11.71 70.27 26.99 26.87 16.96 10.90 7.39 62.77 5.99 6.27 88.29 6.06 6.31 0.00 0.00 0.00
Meditron-7B 33.60 16.33 16.81 57.19 33.96 35.27 26.51 16.48 17.49 29.54 7.52 7.07 21.18 7.35 7.40 12.35 0.29 0.56
@@
MolT5-Small 100.00 0.00 0.00 100.00 0.00 0.00 24.85 9.69 6.54 100.00 0.00 0.00 3.42 0.43 0.09 1.96 0.00 0.00
MolT5-Base 100.00 0.00 0.00 67.45 8.51 8.84 28.00 11.51 6.52 4.17 0.12 0.20 2.20 0.58 0.11 9.70 0.23 0.15
MolT5-Large 100.00 0.00 0.00 69.42 10.29 10.85 15.77 12.28 8.16 18.14 9.52 6.62 8.90 2.28 1.13 4.32 1.16 0.61
Meditron-7B 100.00 0.00 0.00 48.79 11.75 11.05 29.10 20.64 20.64 23.93 12.33 14.34 35.69 19.91 22.65 14.79 9.34 8.98
@@
MolT5-Small 100.00 0.00 0.00 100.00 0.00 0.00 100.00 0.00 0.00 55.49 1.99 1.70 87.44 50.08 49.94 71.86 21.03 24.27
MolT5-Base 100.00 0.00 0.00 100.00 0.00 0.00 100.00 0.00 0.00 58.90 2.25 1.80 94.61 55.16 59.18 45.06 25.49 24.54
@@ 6 Benchmarks
Overall, the dataset proves to be fairly challenging for these naively finetuned models. On captioning, Meditron-7B achieves a maximum overall F-1 score of 16.81 for property identification (Table 3). However, overall it has a much higher precision than recall, indicating the model only labels a molecule with a certain property when having higher confidence. Certain classes of molecules, such as X-icides, are never identified (Table 4). Other classes, such as toxins or electricity, show emergent behavior as model size scales. Interestingly, the models appear to be fairly capable at linking molecules to certain diseases or cancers. We find that, likely due to poor performance on individual properties, only the largest model succeeds on predicting held-out combos, and with poor results. Additionally, we find that the Text2Mol metric, as trained on ChEBI-20, shows poor domain transfer to L+M-24. For the final workshop evaluation, we plan to train an updated Text2Mol model on L+M-24.